OpenStudy (anonymous):

A hawk sitting on top of a tree spots a mouse on the ground 15 meters from the base of the tree. The hawk swoops down toward the mouse at an angle of 30º. What is the distance from the tree branch to the mouse?

OpenStudy (anonymous):

http://prntscr.com/6h5f6t

OpenStudy (welshfella):

you need to use the trig ratio equal to opposite / hypotenuse which one is that?

OpenStudy (anonymous):

sin A = cos(90°−A) ?

OpenStudy (welshfella):

you use the sine because sin = opp/hyp - no need to worry about cos (90 - A)

OpenStudy (welshfella):

just plug in A = 30 and opposite = 15 and solve for h

OpenStudy (welshfella):

sin 39 = 15/h

OpenStudy (welshfella):

* sin 30

OpenStudy (welshfella):

the distance form tree top to the mouse is h

OpenStudy (welshfella):

can you solve for h?